OpenStudy (anonymous):

Simplify the expression: ln(e^sinx^2)+ln(e^cosx^2)

OpenStudy (perl):

ln(e^u)=u

OpenStudy (anonymous):

Then what's left would be sinx^2+cosx^2?

OpenStudy (anonymous):

Yes :)

OpenStudy (perl):

and thats equal to 1, right?

OpenStudy (anonymous):

Who knows, really? :) The problem with these things is that people see them differently... Some people see \[\large \cos x^2 + \sin x ^2\]Others see \[\large (\cos x)^2+(\sin x )^2\] Who's to say? ^_^

OpenStudy (perl):

obviously it means the latter

OpenStudy (anonymous):

Leave that to the asker. Besides, I'm sure if it was the latter, it'd be written like \[\large \cos^2x+\sin^2 x\]

OpenStudy (anonymous):

@peterpan, yeah, that's how it was written.

OpenStudy (anonymous):

Well, then it's settled. Type cos^2 x next time, it will never be misinterpreted. @perl no need to be so mean, I was just making sure.

OpenStudy (anonymous):

Alright, my bad

OpenStudy (perl):

peter, sorry

OpenStudy (perl):

peter, yes the way this person wrote it , it was ambiguous. because he doesnt know better :)

OpenStudy (perl):

but since it said 'simplify' , that would be simpler (1)

OpenStudy (anonymous):

Better safe than sorry. Assuming is dangerous...

OpenStudy (perl):

it depends what youre doing

OpenStudy (perl):

better not to assume if you are running a nuclear reactor. its ok to assume when youre helping with homework

OpenStudy (anonymous):

It's ok, but is it better?

OpenStudy (perl):

better not assume when youre flying a plane,

OpenStudy (perl):

yes, because it can save time

OpenStudy (anonymous):

With a chance at error. Besides, it did cause the asker to rectify the problem and/or settle things instead of leaving us wondering.